OpenStudy (anonymous):

You have a pool that measure 8 yards by 10 yards. You want to build a deck around your pool of uniform width increasing the total area of the pool to 120 square yards. What is the width of the deck

OpenStudy (callisto):

Let x be the width increased. New measure of pool = (x+8) and (x+10). Total area = 120. So the following question can be set: (x+8)(x+10) = 120 Can you solve it?

OpenStudy (anonymous):

i dont rlly kno how..

OpenStudy (callisto):

Can you expand the expression on the left, that is (x+8)(x+10) ?

OpenStudy (callisto):

Hint: break the bracket: (x+8)(x+10) = x(x+10) + 8(x+10) = ...?

OpenStudy (anonymous):

okay i got it now thank you:)

OpenStudy (callisto):

Welcome!